OpenStudy (anonymous):

This question came from a Harvard research study, the answer might surprise you."Bill and Sally’s age totals 28 years. Bill is 20 years older than Sally. How old is Sally? "

OpenStudy (anonymous):

Bill is 24 and Sally is 4

OpenStudy (anonymous):

sally is 4 and bill is 24

OpenStudy (jamesj):

It's a good trap.

OpenStudy (anonymous):

According to the Harvard study if you answered 28 you are intuitive and more likely to believe in God. If you answered 24 you either googled it from my clue or are more likely not to believe things through intuition.

OpenStudy (anonymous):

If you answered Sally was 4 your more likely an athesit according to the Harvard study.

OpenStudy (anonymous):

What an interesting conclusion from the test...weird that it holds true...

OpenStudy (anonymous):

wierder still an almost equal amount of atheisits got it wrong also.

OpenStudy (anonymous):

More likely the correct answer was given by someone who is good at maths/algebra....☺

OpenStudy (anonymous):

no doubt.

OpenStudy (anonymous):

at times sheer reflectiveness is a strength, at other times it is a real weakness-this is the conclusion of the head of Harvards Intuition study.

OpenStudy (anonymous):

well............!!!

OpenStudy (anonymous):

so since I got your attention help me reflect on this answer.Train A and B are traveling in the same direction on parallel tracks.Train A is traveling at 40 MPH and train B is traveling at 48 MPH.Train A passes a train station at 4:15 A.M.. If train B passes the same station at 4:30 a.m. at what time will train B catch up to train A?

OpenStudy (anonymous):

It appears that google can make any one appear wise, for a time.

OpenStudy (anonymous):

Knowledge of maths makes one appear wise ALL THE TIME !!!! Train A passes the station 15 min earlier and in those 15 min it covers 10 miles. After crossing the station at 4:30 am, let train B catch train A after x hours. In those x hours, train A covers distance = x*40 = 40x In the same x hours, train B covers distance = x*48 = 48x But since train B catches up with train A, it means 48x = 40x + 10 48x - 40x = 10 8x = 10 x = 10/8 hours x = (10/8)*60 minutes x = 75 minutes i.e. 1 hour 15 minutes So train B will catch train A at 5:45 am !!!!!!!

OpenStudy (jamesj):

Do you have a link to this alleged Harvard study?